Question 3 Given the data X1 = 1, X2 = 4, X3 = 5, X4 = 8, X5 = 10, evaluate 2 y2 56 0 28 0 e 784 206
Answer #1

Please don't hesitate to give a "thumbs up" for the answer in case the answer has helped you

This is 1^2 + 4^2 + 5^2 + 8^2 + 10^2
= 1+16+25+64 + 100
= 17+89+100
= 206

Answer is 206
Answer #2

ANSWER :


∑ X^2 

= 1^2 + 4^2 + 5^2 + 8^2 + 10^2

= 1 + 16 + 25 + 64 + 100

= 206 : Option D (ANSWER)

answered by: Tulsiram Garg